# New Algebraic Points on Curves

### v1

@@ 4. Counting low degree fields with local restrictions
Let $T$ be a finite set of primes. Write $N_{T}(X)$ for the number of isomorphism classes of cubic fields $K$ satisfying
@@
    $\lvert\Delta_{K}\rvert\leq X$;
@@
Let $T$ be a finite set of primes, with $3\notin T$. Write $N_{T}(X)$ for the number of isomorphism classes of cubic fields $K$ satisfying
@@
    $\lvert\Delta_{K}\rvert\leq X$;
@@
    $\operatorname{ord}_{p}(\Delta_{K})$ is even for all $p\in T$.
@@
Proof of Corollaries 10 and 11.
Theorem 1.3 of [1] expresses the density $C_{T}^{\pm}$ as a product of local densities, which depend on the splitting type of $p$ in $K$ (either totally split, partially split, inert, partially ramified, totally ramified). In Corollary 10 the allowed splitting types for $p\in T$ are totally split, partially split, inert, and the corollary follows immediately from the aforementioned [1, Theorem 1.3].
Next, we note that for $p\neq 3$, the valuation $\operatorname{ord}_{p}(\Delta_{K})$ is either 0, 11 or $2$ according to whether $p$ is unramified in $K$, partially ramified in $K$, or totally ramified in $K$. Thus, condition (ii) is equivalent to saying that for every $p\in T$ the splitting type of $p$ is either totally split, partially split, inert or totally ramified. The result again follows from [1, Theorem 1.3]. ∎
The following is a special case of [7, Theorem 4.1].
Theorem 12 (Ellenberg, Pierce and Wood).
Let $T$ be a finite set of primes. Write $\tilde{N}_{T}(X)$ for the number of isomorphism classes of non-$D_{4}$ quartic fields $K$ satisfying
  • $0<\lvert\Delta_{K}\rvert\leq X$;
  • all $p\in T$ are unramified in $K$.
Then
$\tilde{N}_{T}(X)\;=\;C_{T}\cdot X+O_{T,\epsilon}(X^{23/24+\epsilon})$
where
$C_{T}\;=\;C\cdot\prod_{p\in T}\left(1-\frac{p^{2}+2p+1}{p^{3}+p^{2}+2p+1}\right).$
and $C>0$ is some fixed absolute constant.
Cohen, Diaz y Diaz and Olivier [4] show that the number of $D_{4}$-extensions with absolute discriminant at most $X$ is $\tilde{c}\cdot X$ where $c\approx 0.052326$, but there does not seem to be a count for $D_{4}$-extension with local restrictions.
Finally, the following is a special case of [7, Theorem 5.1].
Theorem 13 (Ellenberg, Pierce and Wood).
Let $T$ be a finite set of primes. Write $N_{T}(X)$ for the number of isomorphism classes of quintic fields $K$ satisfying
  • $0<\lvert\Delta_{K}\rvert\leq X$;
  • all $p\in T$ are unramified in $K$.
Then
$N_{T}(X)\;=\;C_{T}\cdot X+O_{T,\epsilon}(X^{199/200+\epsilon})$
where
$C_{T}\;=\;C\cdot\prod_{p\in T}\left(1-\frac{p^{3}+2p^{2}+2p+1}{p^{4}+p^{3}+2p^{2}+2p+1}\right).$
and $C>0$ is some fixed absolute constant.
@@
$\frac{\sum_{p\in R}\frac{1}{p^{\sigma}}}{-\log(\sigma-1)}\;\leq\;\frac{\sum_{p\in R}\frac{1}{p}}{-\log(\sigma-1)}\rightarrow 0$
since $\sum_{p\in\mathbb{P}}1/p^{\sigma}=-\log(\sigma-1)+O(1)$. This contradicts the assumption that $\delta>0$. ∎
@@
Let $S$ be a set of primes having regular density <1<1.
  1. (a)
    Let $d=2$, 33 or $5$. Then the proportion of degree $d$ number fields with discriminants supported only on $S$ is $0\%$, when ordered by absolute discriminant.
  2. (b)
    Let $d=4$. Then the proportion of non-$D_{4}$ quartic fields with discriminants supported only on $S$ is $0\%$, when ordered by absolute discriminant.
@@
Write $\mathcal{F}(X)$ for the set of quintic fields with absolute discriminant at most $X$, and write $\mathcal{F}_{S}(X)$ for the subset whose discriminants are supported on $S$. To prove the corollary for $d=5$, we want to show that
@@
$0\leq\frac{\#\mathcal{F}_{S}(X)}{\#\mathcal{F}(X)}\leq\frac{N_{p_{1},\dotsc,p_{k}}(X)}{N_{\emptyset}(X)}\rightarrow\prod_{i=1}^{k}\left(1-\frac{p_{i}^{3}+2p_{i}^{2}+2p_{i}+1}{p_{i}^{4}+p_{i}^{3}+2p_{i}^{2}+2p_{i}+1}\right)$
as $X\rightarrow\infty$, where the limit follows from Theorem 13. From Lemma 14, by choosing $p_{1},\dotsc,p_{k}$ be a suitable (finite) subset of $R$, we can ensure that the product on the right hand-side is as small as we like. This completes the proof of (a) for $d=5$. The other cases are similar. ∎
@@
Let $F(T,Y)\in\mathbb{Q}(T)[Y]$ has degree $3$ in $Y$, and be irreducible over $\mathbb{Q}(T)$, and $\Theta$ be the set of $t\in\mathbb{Q}$ such that $F(t,Y)$ either has degree <n<n or is reducible (this set is necessarily thin by Hilbert’s irreducibility theorem). Let $\Delta_{Y}(F)\in\mathbb{Q}(T)$ be the discriminant of $F$ with respect to $Y$, and write this as
@@
where $g\in\mathbb{Q}[T]$ is a squarefree polynomial and $h(T)\in\mathbb{Q}(T)$ is a rational function. Suppose $g$ has even degree. Let $G=\operatorname{Gal}(g)$. Suppose that there is an element $\sigma\in\operatorname{Gal}(g)$ acting freely on the roots of $g$. Let $\mathcal{F}^{\prime}$ be the set of cubic fields we obtain from $F(t,Y)=0$ with $t\in\mathbb{Q}-\Theta$. Then $100\%$ of cubic fields, ordered by discriminant, do not belong to the family $\mathcal{F}^{\prime}$.
@@
By Lemma 8, there is a set $S$ of primes having positive regular density, such that for all $t\in\mathbb{Q}-\Theta$ and all $p\in S$, the valuation $\operatorname{ord}_{p}(\Delta(\mathcal{O}_{K_{t}}))$ is even, for any $K_{t}\in\mathcal{F}^{\prime}$. The desired result follows from Corollary 16. ∎
@@ 5. Proof of Theorem 4
Let $L$ be a number field. Nagell [19] calls a unit $\varepsilon\in\mathcal{O}_{L}^{\times}$ exceptional if $1-\varepsilon\in\mathcal{O}_{L}^{\times}$. The number field $L$ is called exceptional if it possesses an exceptional unit. Thus $\varepsilon$ is exceptional if and only if $(\varepsilon,1-\varepsilon)$ is a solution to the unit equation (1), and $L$ is exceptional if and only if the unit equation has solutions in $L$. Nagell [17], [18], has determined all exceptional number fields where the unit rank is 0 or $1$. Nagell finds [18, Section 2] that the only exceptional quadratic fields are $\mathbb{Q}(\sqrt{5})$ and $\mathbb{Q}(\sqrt{-3})$ which contain exceptional units $(1+\sqrt{5})/2$ and $(1-\sqrt{-3})/2$ respectively, and the only exceptional complex cubic fields are the ones with discriminants 23-23 and 31-31. He also showed [18, Sections 3–5] that the only exceptional real cubic fields are of the form $\mathbb{Q}(\varepsilon)$ where $\varepsilon$ is a root of
@@
It follows that if a prime $p$ divides the discriminant of $\mathcal{O}_{\mathbb{Q}(\varepsilon)}$ then it divides $t^{4}+6t^{3}+7t^{2}-6t-31$ and hence the polynomial $F(X)$ has a root modulo $p$. Write $S$ for the set of primes $p$ such that $F$ has a root modulo $p$. The Galois group of $F$ is $D_{4}$ and an easy application of the Chebotarev Density Theorem shows that $S$ has Frobenian (and hence regular) density $3/8$.
Write $\mathcal{F}_{S}$ for the set of cubic fields with discriminants supported on $S$. By the above, the exceptional fields with Galois group $S_{3}$ all belong to $\mathcal{F}_{S}$. The latter has density 0 by Corollary 15. This completes the proof.
@@ 6. Proof of Theorem 5
Using Magma [2] we computed the Riemann–Roch space $L(D_{a,b,c})$ for the $32$ possible combinations $(a,b,c)$. For all but $4$ of these, the Riemann–Roch dimension is $1$, and therefore $D=D_{a,b,c}$. For the remaining $4$, namely $(a,b,c)=(0,0,0)$, $(0,1,0)$, $(0,1,1)$, $(1,2,1)$, the Riemann–Roch dimension is $2$, and therefore there are infinitely many effective degree $3$ divisors linearly equivalent to $D_{a,b,c}$. Consider the case $(a,b,c)=(0,0,0)$. Then $D_{(0,0,0)}=3P_{1}$, and a basis for $L(D_{(0,0,0)})$ is given by $1$, $\varphi=y/x$, where these are considered as elements of the function field $\mathbb{Q}(X_{0}(64))$. Thus

### v2

@@ 5. Proof of Theorem 6
Let $T$ be a finite set of primes. Write $N_{T}(X)$ for the number of isomorphism classes of cubic fields $L$ satisfying
@@
    $\lvert\Delta_{L}\rvert\leq X$;
@@
Let $T$ be a finite set of primes, with $2$, $3\notin T$. Write $N_{T}(X)$ for the number of isomorphism classes of cubic fields $L$ satisfying
@@
    $\lvert\Delta_{L}\rvert\leq X$;
@@
    $\operatorname{ord}_{p}(\Delta_{L})$ is even for all $p\in T$.
@@
Proof of Corollaries 13 and 14.
Theorem 1.3 of [1] expresses the density $C_{T}$ as a product of local densities, which depend on the splitting type of $p$ in $L$ (either totally split, partially split, inert, partially ramified, totally ramified). In Corollary 13 the allowed splitting types for $p\in T$ are totally split, partially split, inert, and the corollary follows immediately from the aforementioned [1, Theorem 1.3].
We turn to the proof of Corollary 14. For this we note that for $p\neq 2$, $3$, the valuation $\operatorname{ord}_{p}(\Delta_{L})$ is either 0, 11 or $2$ according to whether $p$ is unramified in $L$, partially ramified in $L$, or totally ramified in $L$; this follows immediately from Lemma 15 below. Thus, condition (ii) of Corollary 14 is equivalent to saying that for every $p\in T$ the splitting type of $p$ is either totally split, partially split, inert or totally ramified. The result again follows from [1, Theorem 1.3]. ∎
The following result is well-known. We include the proof for the convenience of the reader.
Lemma 15.
Let $L$ be a number field of degree $n$. Let $p>n$ be a rational prime and write
$p\mathcal{O}_{L}=\mathfrak{P}_{1}^{e_{1}}\cdots\mathfrak{P}_{r}^{e_{r}}$
where the $\mathfrak{P}_{i}$ are distinct prime ideals. Let $f_{i}$ be the inertial degree of $\mathfrak{P}_{i}$. Then
$\operatorname{ord}_{p}(\Delta_{L})=\sum_{i=1}^{r}f_{i}(e_{i}-1).$
Proof.
The condition $p>n$ ensures that $p\nmid e_{i}$, so the primes $\mathfrak{P}_{i}$ are tamely ramified in $L$. Write $\mathcal{D}_{L}$ for the different ideal of $L$. Then $\operatorname{ord}_{\mathfrak{P}_{i}}(\mathcal{D}_{L})=e_{i}-1$ (Theorem III.2.6 of [26]). The lemma follows as $\operatorname{Norm}(\mathcal{D}_{L})=\lvert\Delta_{L}\rvert$ (Theorem III.2.9 of [26]). ∎
@@
$\frac{\sum_{p\in R}\frac{1}{p^{\sigma}}}{-\log(\sigma-1)}\;\leq\;\frac{\sum_{p\in R}\frac{1}{p}}{-\log(\sigma-1)}\rightarrow 0.$
This contradicts the assumption that $\delta>0$. ∎
@@
Let $S$ be a set of primes having regular density <1<1. Then the proportion of cubic number fields with discriminants supported only on $S$ is $0\%$, when ordered by absolute discriminant.
@@
Write $\mathcal{F}(X)$ for the set of cubic fields with absolute discriminant at most $X$, and write $\mathcal{F}_{S}(X)$ for the subset whose discriminants are supported on $S$. We want to show that
@@
$0\leq\frac{\#\mathcal{F}_{S}(X)}{\#\mathcal{F}(X)}\leq\frac{N_{p_{1},\dotsc,p_{k}}(X)}{N_{\emptyset}(X)}\rightarrow\prod_{i=1}^{k}\left(1-\frac{p_{i}+1}{p_{i}^{2}+p_{i}+1}\right)$
as $X\rightarrow\infty$, where the limit follows from Corollary 13. From Lemma 16, by choosing $p_{1},\dotsc,p_{k}$ be a suitable (finite) subset of $R$, we can ensure that the product on the right hand-side is as small as we like. This completes the proof. ∎
@@
Let $F(T,Y)\in\mathbb{Q}[T,Y]$ be irreducible and have degree $3$ in $Y$. Let $\Theta$ be the set of $t\in\mathbb{Q}$ such that $F(t,Y)$ either has degree <3<3 or is reducible. Let $\Delta_{Y}(F)\in\mathbb{Q}[T]$ be the discriminant of $F$ with respect to $Y$, and write this as
@@
where $g$, $h\in\mathbb{Q}[T]$ with $g$ squarefree. Suppose $g$ has even degree. Let $G=\operatorname{Gal}(g)$. Suppose that there is an element $\sigma\in\operatorname{Gal}(g)$ acting freely on the roots of $g$. Let $\mathcal{F}^{\prime}$ be the set of cubic fields we obtain from $F(t,Y)=0$ with $t\in\mathbb{Q}-\Theta$. Then $100\%$ of cubic fields, ordered by discriminant, do not belong to the family $\mathcal{F}^{\prime}$.
@@
By Lemma 12, there is a set $S$ of primes having positive regular density, such that for all $t\in\mathbb{Q}-\Theta$ and all $p\in S$, the valuation $\operatorname{ord}_{p}(\Delta(\mathcal{O}_{L_{t}}))$ is even, for any $L_{t}\in\mathcal{F}^{\prime}$. The desired result follows from Corollary 18. ∎
Proof of Theorem 6.
We apply the proof of Theorem 9 with $n=3$. We deduce that all but finitely many cubic points on $C$ arise as fibres of the map $C\rightarrow\mathbb{P}^{1}$ given on the affine model $F(T,Y)=0$ by $(T,Y)\mapsto T$; the assumption $\operatorname{genus}(C)\geq 8$ ensures that $C$ satisfies condition (3). Theorem 6 follows immediately from Corollary 19. ∎
@@ 6. Proof of Theorem 7
Let $L$ be a number field. Nagell [24] calls a unit $\varepsilon\in\mathcal{O}_{L}^{\times}$ exceptional if $1-\varepsilon\in\mathcal{O}_{L}^{\times}$. The number field $L$ is called exceptional if it possesses an exceptional unit. Thus $\varepsilon$ is exceptional if and only if $(\varepsilon,1-\varepsilon)$ is a solution to the unit equation (1), and $L$ is exceptional if and only if the unit equation has solutions in $L$. Nagell [22], [23], has determined all exceptional number fields where the unit rank is 0 or $1$. Nagell finds [23, Section 2] that the only exceptional quadratic fields are $\mathbb{Q}(\sqrt{5})$ and $\mathbb{Q}(\sqrt{-3})$ which contain exceptional units $(1+\sqrt{5})/2$ and $(1+\sqrt{-3})/2$ respectively, and the only exceptional complex cubic fields are the ones with discriminants 23-23 and 31-31. He also showed [23, Sections 3–5] that the only exceptional real cubic fields are of the form $\mathbb{Q}(\varepsilon)$ where $\varepsilon$ is a root of
@@
It follows that if a prime $p$ divides the discriminant of $\mathcal{O}_{\mathbb{Q}(\varepsilon)}$ then it divides $F(t)$ and hence the polynomial $F$ has a root modulo $p$. Write $S$ for the set of primes $p$ such that $F$ has a root modulo $p$. The Galois group of $F$ is $D_{4}$ and an easy application of the Chebotarev Density Theorem shows that $S$ has Frobenian (and hence regular) density $3/8$.
Write $\mathcal{F}_{S}$ for the set of cubic fields with discriminants supported on $S$. By the above, the real exceptional fields with Galois group $S_{3}$ all belong to $\mathcal{F}_{S}$. The latter has density 0 by Corollary 17. This completes the proof. ∎
@@ 7. Proof of Theorem 8
Using Magma [2] we computed the Riemann–Roch space $L(D_{a,b,c})$ for the $32$ possible combinations $(a,b,c)$. For all but $4$ of these, the Riemann–Roch dimension is $1$, and therefore there is precisely one effective $D$ linearly equivalent to $D_{a,b,c}$. For the remaining $4$, namely $(a,b,c)=(0,0,0)$, $(0,1,0)$, $(0,1,1)$, $(1,2,1)$, the Riemann–Roch dimension is $2$, and therefore there are infinitely many effective degree $3$ divisors linearly equivalent to $D_{a,b,c}$. Consider the case $(a,b,c)=(0,0,0)$. Then $D_{(0,0,0)}=3P_{1}$, and a basis for $L(D_{(0,0,0)})$ is given by $1$, $\varphi\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}y/x$, where these are considered as elements of the function field $\mathbb{Q}(X_{0}(64))$. Thus
@@
8. Suggested Future Projects
We invite the reader to help gather further evidence towards our Conjecture 2. Here are some projects that we believe are ‘low-hanging fruit’:
  1. (I)
    Theorem 9 gives an upper bound for the number of discriminants of degree $n$ primitive number fields coming from degree $n$ new points on a curve $C$. It assumed that $C$ has a plane model of the form $F(T,Y)=0$ where $F$ has degree $n$ in $Y$. This assumption is equivalent to the existence of a degree $n$ morphism $\psi:C\rightarrow\mathbb{P}^{1}$. We expect that it should be possible to prove a similar theorem where there is a degree $n$ morphism $\psi:C\rightarrow E$ to an elliptic curve $E$. Here, instead of Vojta’s Theorem (recounted in Section 4), it would be useful to apply a more general theorem due to Song and Tucker [31, Proposition 2.3].
  2. (II)
    The proof of Theorem 6 makes use of a theorem of Bhargava, Taniguchi and Thorne [1, Theorem 1.3] that counts cubic fields, ordered by discriminant, with local specifications. Similar theorems for degrees $4$ and $5$ have been established by Ellenberg, Pierce and Wood [9, Theorems 4.1 and 5.1]. As far as we can tell, to prove the analogue of Theorem 6 for degrees $4$ and $5$ would require a refinement of the aforementioned theorems of Ellenberg, Pierce and Wood, that consider densities of different possible ramification types for primes.
  3. (III)
    Corollary 5 and Theorem 8 establish Conjecture 3 in degrees $2$, 33 for certain $X_{0}(N)$. It should be possible to establish many similar results for other modular curves appearing in the $\mathrm{L}$-Functions and Modular Forms database [33].